# Leveraging Symmetry to Accelerate Learning of Trajectory Tracking Controllers for Free-Flying Robotic Systems

### v2

@@ Abstract
Tracking controllers enable robotic systems to accurately follow planned reference trajectories. In particular, reinforcement learning (RL) has shown promise in the synthesis of controllers for systems with complex dynamics and modest online compute budgets. However, the poor sample efficiency of RL and the challenges of reward design make training slow and sometimes unstable, especially for high-dimensional systems. In this work, we leverage the inherent Lie group symmetries of robotic systems with a floating base to mitigate these challenges when learning tracking controllers. We model a general tracking problem as a Markov decision process (MDP) that captures the evolution of both the physical and reference states. Next, we prove that symmetry in the underlying dynamics and running costs leads to an MDP homomorphism, a mapping that allows a policy trained on a lower-dimensional “quotient” MDP to be lifted to an optimal tracking controller for the original system. We compare this symmetry-informed approach to an unstructured baseline, using Proximal Policy Optimization (PPO) to learn tracking controllers for three systems: the Particle (a forced point mass), the Astrobee (a fully-actuated space robot), and the Quadrotor (an underactuated system). Results show that a symmetry-aware approach both accelerates training and reduces tracking error after the same number of training steps.
@@ II-A Homomorphisms of Markov Decision Processes
for all ${s\in\mathcal{S}}$, ${a\in\mathcal{A}}$, and ${\widetilde{B}\in\mathsf{B}({\widetilde{S}}})$. Given a continuous MDP homomorphism $(p,h)$, a policy $\pi$ for $\widetilde{\mathcal{M}}$, and a policy $\widetilde{\pi}$ for $\mathcal{M}$, $\pi$ is called a lift of $\widetilde{\pi}$ if for all ${s\in\mathcal{S}}$ and ${A\in\mathsf{B}(\mathcal{A})}$,
@@ III-A Modeling a Tracking Control Problem as an MDP
We model the tracking task for reference trajectories that are unknown a priori in the following manner.
@@ III-B Symmetries of Tracking Control MDPs
where we have substituted in (17a) and simplified using (15). Thus, (5a) holds. Considering now the transitions, we note that (12) can also be written using the “product measure” as
@@
We then apply (17a) to (21) to compute
@@
Considering the Lie groups ${\mathcal{K}:=T\mathbb{R}^{3}}$ (with the group operation inherited from its identification with ${\mathbb{R}^{3}\times\mathbb{R}^{3}}$) and ${\mathcal{H}:=\mathbb{R}^{3}}$, we let an $\mathcal{K}$-action on ${\mathcal{S}=T\mathbb{R}^{3}}$ and an $\mathcal{H}$-action on ${\mathcal{A}=\mathbb{R}^{3}}$ be given by the left action of the groups on themselves, i.e.,
@@ VI Experiments
We use a custom implementation of PPO [29] (see code for details), with the same hyperparameters across all variants of each environment. During training, the reference actions are sampled from a stationary distribution (as in Def. 5), but we evaluate zero-shot on pre-planned (dynamically feasible) reference trajectories. Fig. 1 and Table I report total reward (during training) and average tracking error (during evaluation), starting the system from a randomized initial state.
(a) Particle
(b) Astrobee
(c) Quadrotor
Figure 1: Reward during training and tracking error components during evaluation for the Particle, Astrobee, and Quadrotor, with translational errors as solid lines and rotational errors (when applicable) as dashed lines.
@@
Environment $\mathcal{G}$ $r\,\mathrm{[m]}$ $v\,\mathrm{[m/s]}$ $R\,\mathrm{[rad]}$ $\omega\,\mathrm{[rad/s]}$
  Particle Baseline 2.50$\pm$0.25 0.93$\pm$0.11 - -
$\mathbb{R}^{3}$ 0.12$\pm$0.4 0.06$\pm$0.23 - -
$T\mathbb{R}^{3}$ 0.11$\pm$0.41 0.04$\pm$0.19 - -
$T\mathbb{R}^{3}\times\mathbb{R}^{3}$ 0.09$\pm$0.42 0.04$\pm$0.19 - -
Astrobee Baseline 0.23$\pm$0.01 0.17$\pm$0.01 0.70$\pm$0.03 1.91$\pm$0.08
$SE(3)$ 0.10$\pm$0.01 0.03$\pm$0.04 0.41$\pm$0.04 1.51$\pm$0.07
Quadrotor Baseline 0.91$\pm$0.62 0.97$\pm$0.78 0.51$\pm$0.19 0.27$\pm$0.21
$SE(2)\times\mathbb{R}$ 0.25$\pm$0.30 0.14$\pm$0.26 0.29$\pm$0.02 0.04$\pm$0.07
We report the mean and standard deviation (over ${n=20}$ training seeds) of the policy’s RMS tracking error (on a dataset of $m=20$ trajectories).
@@ VII Discussion
Fig. 1 shows a clear trend across the board: greater symmetry exploitation leads to improved sample efficiency. The tracking error evaluation shown in Table I and Fig. 1 follows a similar trend. For the Particle, the vast majority of this benefit is achieved by reduction of the translational symmetry, although incorporating the velocity and force symmetries yields modest additional gains. This seems consistent with the large improvement we see for the Astrobee and Quadrotor after reduction by (a subgroup of) $SE(3)$. Careful reward engineering or hyperparameter tuning might improve performance (especially for the baseline, which currently fails to learn effectively), but we instead focus on analyzing the benefit of exploiting symmetry for a fixed reward. Nonetheless, any reward depending only on the reduced state $\tilde{s}=p(s)$ would preserve the symmetry.

### v3

@@ Abstract
Tracking controllers enable robotic systems to accurately follow planned reference trajectories. In particular, reinforcement learning (RL) has shown promise in the synthesis of controllers for systems with complex dynamics and modest online compute budgets. However, the poor sample efficiency of RL and the challenges of reward design make training slow and sometimes unstable, especially for high-dimensional systems. In this work, we leverage the inherent Lie group symmetries of robotic systems with a floating base to mitigate these challenges when learning tracking controllers. We model a general tracking problem as a Markov decision process (MDP) that captures the evolution of both the physical and reference states. Next, we prove that symmetry in the underlying dynamics and running costs leads to an MDP homomorphism, a mapping that allows a policy trained on a lower-dimensional “quotient” MDP to be lifted to an optimal tracking controller for the original system. We compare this symmetry-informed approach to an unstructured baseline, using Proximal Policy Optimization (PPO) to learn tracking controllers for three systems: the Particle (a forced point mass), the Astrobee (a fully-actuated space robot), and the Quadrotor (an underactuated system). Results show that a symmetry-aware approach both accelerates training and reduces tracking error at convergence.
@@ II-A Homomorphisms of Markov Decision Processes
for all ${s\in\mathcal{S}}$, ${a\in\mathcal{A}}$, and ${\widetilde{B}\in\mathsf{B}({\widetilde{S}}})$. Given a continuous MDP homomorphism $(p,h)$, a policy $\widetilde{\pi}$ for $\widetilde{\mathcal{M}}$, and a policy $\pi$ for $\mathcal{M}$, $\pi$ is called a lift of $\widetilde{\pi}$ if for all ${s\in\mathcal{S}}$ and ${A\in\mathsf{B}(\mathcal{A})}$,
@@ III-A Modeling a Tracking Control Problem as an MDP
We choose to model the trajectory tracking task for a priori unknown reference trajectories in the following manner.
@@ III-B Symmetries of Tracking Control MDPs
where we have substituted in (17) and simplified using (15). Thus, (5a) holds. Considering now the transitions, we note that (12) can also be written using the “product measure” as
@@
We then apply (17) to (21) to compute
@@
Considering the Lie groups ${\mathcal{K}:=T\mathbb{R}^{3}}$ (with the obvious group operation inherited from its identification with ${\mathbb{R}^{3}\times\mathbb{R}^{3}}$) and ${\mathcal{H}:=\mathbb{R}^{3}}$, we let a $\mathcal{K}$-action on ${\mathcal{S}=T\mathbb{R}^{3}}$ and an $\mathcal{H}$-action on ${\mathcal{A}=\mathbb{R}^{3}}$ be given by the left action of the groups on themselves, i.e.,
@@ VI Experiments
We use a custom implementation of PPO [29] (see code for details), with the same hyperparameters across all variants of each environment. During training, the reference actions are sampled from a stationary distribution (as in Def. 5), but we evaluate zero-shot on pre-planned (dynamically feasible) reference trajectories. Fig. 1 and Table I report total reward (during training) and average tracking error (during evaluation) when starting from a randomly sampled initial state.
(a) Training and evaluation for Particle.
(b) Training and evaluation for Astrobee.
(c) Training and evaluation for Quadrotor.
Figure 1: Mean reward during training (for 10 training seeds) and, for the best-performing seed, mean tracking error during evaluation (for 20 trajectories), with translational errors as solid lines and rotational errors as dashed lines.
@@
Environment $\mathcal{G}$ $r\,\mathrm{[cm]}$ $v\,\mathrm{[cm/s]}$ $R\,\mathrm{[rad]}$ $\omega\,\mathrm{[rad/s]}$
  Particle Baseline 298.3$\pm$1.2 103.0$\pm$1.1 - -
$\mathbb{R}^{3}$ 010.0$\pm$0.1 004.8$\pm$0.1 - -
$T\mathbb{R}^{3}$ 009.8$\pm$0.2 $0$
$T\mathbb{R}^{3}\times\mathbb{R}^{3}$ 008.3$\pm$0.1 003.9$\pm$0.2 - -
Astrobee Baseline 010.4$\pm$1.0 006.0$\pm$1.0 0.75$\pm$0.01 0.24$\pm$0.02
$SE(3)$ 001.3$\pm$2.7 003.3$\pm$3.1 0.37$\pm$0.04 0.16$\pm$0.06
Quadrotor Baseline 066.3$\pm$3.2 041.3$\pm$2.5 0.48$\pm$0.02 0.25$\pm$0.02
${SE(2)\times\mathbb{R}}$ 028.4$\pm$4.1 019.6$\pm$3.1 0.25$\pm$0.05 0.12$\pm$0.01
We report the mean and standard deviation (for 1010{10}10 training seeds) of the trained policy’s RMS tracking error (for a dataset of 2020{20}20 trajectories).
@@ VII Discussion
Fig. 1 shows a clear trend across the board: greater symmetry exploitation leads to improved sample efficiency. The tracking error evaluation shown in Table I and Fig. 1 follows a similar trend. For the Particle, the vast majority of this benefit is achieved by reduction of the translational symmetry, although incorporating the velocity and force symmetries yields modest additional gains. This seems consistent with the large improvement we see for the Astrobee and Quadrotor after reduction by (a subgroup of) $SE(3)$. Careful reward engineering or hyperparameter tuning might improve performance (especially for the baseline, which currently struggles to learn effectively), but we instead focus on analyzing the benefit of exploiting symmetry for a fixed reward. Nonetheless, any reward depending only on the reduced state $\tilde{s}=p(s)$ would preserve the symmetry.